OpenStudy (anonymous):

Fair tickets for 2 adults and 3 children cost $34. An adult ticket costs $2 more than a child ticket. What is the price of an adult ticket?

OpenStudy (abdela25):

simple easy division lets go over it step by step

OpenStudy (anonymous):

I don't think that's the right answer, @Onister

OpenStudy (abdela25):

this is how its done lets make it simpler by replacing numbers with variables 2A+3c=24 A=C+2

OpenStudy (abdela25):

do you get it

OpenStudy (abdela25):

sorry for the typo i meant 2A+3C=34 not 24

OpenStudy (anonymous):

Yes I get it... So like wouldn't it be like 2(c + 2) + 3c =34 ...?

OpenStudy (abdela25):

yeah kind of

OpenStudy (abdela25):

substituting 2(c+2)+3c=34 - just like you said

OpenStudy (abdela25):

2c+4+3c=34 5c=30 which equals C=6 a childs ticket is $6 A=C+2 or now A=6+2=8 an Adult ticket costs $8 and a childs ticket cost $6

OpenStudy (abdela25):

easy enough? right?

OpenStudy (anonymous):

Haha yeah, gahhh I just usually over think the problems which make it harder for me.

OpenStudy (abdela25):

yeah ikr thats me all the time lol

OpenStudy (abdela25):

its not that hard if you think about it same with my questions before and btw thanks for the medal

OpenStudy (anonymous):

Yeah, my dad is always yelling at me cuz I over think it all when actually all problems are simple :c And yeah no problem c:

OpenStudy (abdela25):

yeah lol same here

OpenStudy (anonymous):

Haha yeah. But thanks again for your help c:

OpenStudy (abdela25):

ya np your welcome see ya later

OpenStudy (anonymous):

Okay c: